OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. (y + 5) + y = 6 B. x + (y + 5) = 6 C. x + y = y + 5 D. x = (x – 6) + 5

OpenStudy (danjs):

The second equation is solved for x already, put y+5 into the first equation where you see the x,

OpenStudy (danjs):

x + y = 6 x = y + 5

OpenStudy (texaschic101):

exactly ^^ just sub in y + 5 in for x back into the 1st equation

OpenStudy (anonymous):

its A

OpenStudy (danjs):

yea

OpenStudy (anonymous):

im pretty sure it is

OpenStudy (texaschic101):

good job :)

OpenStudy (anonymous):

can you guys help me with another one?

OpenStudy (danjs):

If one wasnt already solved for a variable, like x = y + 5 you would have to do an extra step and solve for a variable in one of the equations first

OpenStudy (danjs):

sure post er up

OpenStudy (danjs):

For example, this is the same question here x + y = 6 x - y = 5 But you have to isolate a variable first, they already did that for you. x + y = 6 x = 5 + y

OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? 5x+_y=2 _x+3y=8

OpenStudy (danjs):

Inconsistent means there are no solutions, where the two lines intersect. They are parallel lines. 5x + _y = 2 _x + 3y = 8 When 2 lines are parallel, the coefficients in front of the X and the Y in those 2 equations will be a multiple of one another.

OpenStudy (anonymous):

Mhm

OpenStudy (danjs):

if you say the first one is 5x + 1y = 2 _x + 3y = 8 if you let the first equation have a 1 on the Y variable, you can see moving to the second equation, the 3y, is 3 times the first equations y

OpenStudy (danjs):

The second equation will be 3 times the first on the x and y variables if you let the first equations y be 1. so the x variable in the second equation will be 3 times the 5x, which is 15x

OpenStudy (danjs):

5x + 1y = 2 15x + 3y = 8 Now the second equations left side is 3 times the first

OpenStudy (anonymous):

then what about the other value

OpenStudy (danjs):

the right side?

OpenStudy (anonymous):

ya

OpenStudy (danjs):

Parallel lines, Inconsistent, zero solutions 5x + 1y = 2 15x + 3y = 8 If the right side was also the same multiple of 3 like the left side is, for example 5x + 1y = 2 15x + 3y = 6 The lines are the same exact line, you can divide the second equation by 3 and get the first back again. Here there are infinite solutions, dependent, overlapping lines

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

k let me see

OpenStudy (danjs):

For example if you have... x + y = 1 2x + 2y = 2 those are the same line

OpenStudy (danjs):

if you have x + y = 1 2x + 2y = 34 those are parallel lines, the left is multiplied by 2, but the right is different than that.

OpenStudy (anonymous):

is it 10?

OpenStudy (danjs):

what do you mean?

OpenStudy (anonymous):

the solution left

OpenStudy (anonymous):

10y?